Let C be the curve of intersection of the parabolic cylinder y=x^2/2 and z=xy/3. What is the length of this curve between the origin and the point (6,18,36)? You'll want to parameterize by t as simply as possible (e.g. x(t) = t ).
1
Expert's answer
2020-09-23T16:09:37-0400
"Solution"

"y= \frac{1}{2}x^2 .............eq 1"

"z= \frac{1}{3}xy.............eq2"

Change equation 1 into a parametric form replacing x with t i.e.


"when, x=t, y= \frac{1}{2}t^2"

Solving equation 2 for z in terms of t, we have;


"z= \frac{1}{3}(x.y) = \frac{1}{3} (t. \frac{1}{2}t^2 ) = \frac{1}{6}t^3"


plugging in the x, y, and z coordinates into a vector equation for the curve r(t), we have;


"r(t) = < t, \frac{1}{2}t^2 , \frac{1}{6}t^3 >"


Now calculating the 1st derivative of r(t)


"r'(t) = < 1, t, \frac{1}{2}t^2 >"


Calculating and simplifying the magnitude of r'(t), we have;


"\begin{vmatrix}
 r'(t) \\
 
\end{vmatrix} = \sqrt{\frac{1}{4}t^4 + t^2 +1}"


"=" "\frac{1}{2} \sqrt{ t^4 + 4 t^2 +4}"


"=" "\frac{1}{2} \sqrt{( t^2 +2)^2}"


"= \frac{1}{2}t^2 +1"


Now finding the range of t along the curve between the origin and the point (6, 18, 36) i.e.


At the origin: "(0,0,0) \u2192 t=0"


At the point: "(6,18, 36) t \u2192 6"


Therefore, the range of t is; "0 \u2264 t \u2264 6"


Now length of the curve, C is obtained by taking the integral for the arclength from 0 to 6, i.e.



"C=\int _{0}^6 (\frac{1}{2}t^2 + 1 ) dt"


"C = (\frac{1}{6}t^3) | _{0}^6 = 42"


Therefore the length of the curve is 42 Units